# Improving Long-Range Navigation with Spatially-Enhanced Recurrent Memory via End-to-End Reinforcement Learning

### v1

@@ 7 Conclusion
In this study, we identify and address a limitation of existing recurrent neural network architectures: their inadequate capability for spatial memorization and the transformation of observations from varying perspectives, which are essential for effective long-range navigation using egocentric perception. To overcome this limitation, we propose Spatially-Enhanced Recurrent Units (SRUs), which integrate an implicit spatial transformation operation into standard GRU and LSTM structures. To enhance navigation performance, SRUs are integrated into a novel attention-based architecture, trained end-to-end via reinforcement learning, leveraging a single forward-facing depth camera for long-range navigation tasks. Our research further highlights the importance of regularization strategies in end-to-end reinforcement learning frameworks. Techniques such as temporally consistent dropout and deep mutual learning are crucial for fully leveraging SRUs’ potential and preventing early overfitting. Experiments demonstrate SRUs’ superior navigation performance compared to standard LSTM and GRU models. Moreover, we compare our implicit recurrent memory-based approach with a state-of-the-art baseline that utilizes explicit mapping and historical paths. Our findings illustrate the superior effectiveness of recurrent memory structures for long-range navigation tasks. Additionally, through ablation studies, we demonstrate the role of specific design choices, particularly the spatial attention mechanism, in enhancing overall navigation performance. Lastly, we analyze and address the challenge of sim-to-real transfer for stereo depth perception by integrating large-scale pretraining. This approach enables successful zero-shot transfer and robust generalization across diverse real-world environments, including indoor, outdoor, and forest scenarios, underscoring the practical applicability and effectiveness of our proposed methodology.
@@ Appendix A Training Details for Spatial-Temporal Memorization Task
  • The ego-motion $M^{t-1}_{t}$ for each step is uniformly sampled, with translation in the range $[-2,2]$ meters and orientation in the range $[-\pi,\pi]$ radians.
  • The observed landmark coordinates $\{l^{i}_{T},i=1,2,\ldots,T\}$ are uniformly sampled within the range $[-5,5]$ meters relative to the observation frame, and the binary categorical labels $\{c^{i},i=1,2,\ldots,T\}$ are uniformly sampled from the set $\{0,1\}$.
@@ Appendix B Parallelizable Stereo Depth Perception Noise Implementation
0:  $depth\in\mathbb{R}^{B\times H\times W}$
  $(K_{mean},K_{sub})\leftarrow$ ComputeKernels($filt\_size$)
@@
  $filtered\_disp\leftarrow$ FilterDisp($disp$, $K_{mean}$, $K_{sub}$)
@@
  return  $filtered\_depth$
@@
  function FilterDisp($disp$, $K_{mean}$, $K_{sub}$)
@@
  $filled\leftarrow\text{if }den>0\text{ then }\frac{num}{den}\text{ else }masked$
  return  $filled$
@@ Appendix C Training Details for Navigation with Reinforcement Learning
Linear Velocity ($v_{t}$) ±$\pm 0.2$ m/s
Angular Velocity ($\omega_{t}$) ±$\pm 0.1$ rad/s
Projected Gravity ($n_{t}$) ±$\pm 0.1$
Goal Position ($p_{t}$) ±$\pm 0.5$ m, ±$\pm 0.1$ rad

### v2

@@ 7 Conclusion
In this study, we identify and address a limitation of existing recurrent neural network architectures in the context of navigation: while RNNs excel at modeling temporal sequences, they are not inherently designed for spatial memorization or transforming observations from varying perspectives. This limitation makes them less effective in building the spatial representations required for mapless navigation using egocentric perception. To address this, we propose Spatially-Enhanced Recurrent Units (SRUs), which integrate an implicit spatial transformation operation into standard GRU and LSTM structures. These SRUs are incorporated into a novel attention-based architecture, trained end-to-end via reinforcement learning, achieving long-horizon navigation tasks with a single forward-facing depth camera. Our research further highlights the importance of regularization strategies in end-to-end reinforcement learning frameworks. Techniques such as temporally consistent dropout and deep mutual learning are crucial for fully leveraging SRUs’ potential and preventing early overfitting. Experiments demonstrate SRUs’ superior navigation performance compared to standard LSTM and GRU models. Moreover, we compare our implicit recurrent memory-based approach with a state-of-the-art baseline that utilizes explicit mapping and historical paths. Our findings illustrate the superior effectiveness of recurrent memory structures for long-range mapless navigation tasks. Additionally, through ablation studies, we demonstrate the role of specific design choices, particularly the spatial attention mechanism, in enhancing overall navigation performance. Lastly, we analyze and address the challenge of sim-to-real transfer for stereo depth perception by integrating large-scale pretraining. This approach enables successful zero-shot transfer and robust generalization across diverse real-world environments, including indoor, outdoor, and forest scenarios, underscoring the practical applicability and effectiveness of our proposed methodology.
@@ Appendix A Training Details for Spatial-Temporal Memorization Task
  • The ego-motion $M^{t-1}_{t}$ for each step is uniformly sampled, with translation in the range [2,2][-2,2] meters and orientation in the range $\pi,\pi]$ radians.
  • The observed landmark coordinates $\{l^{i}_{T},i=1,2,\ldots,T\}$ are uniformly sampled within the range [5,5][-5,5] meters relative to the observation frame, and the binary categorical labels $\{c^{i},i=1,2,\ldots,T\}$ are uniformly sampled from the set $\{0,1\}$.
@@ Appendix B Parallelizable Stereo Depth Perception Noise Implementation
[…]
$(K_{mean},K_{sub})\leftarrow$ ComputeKernels(filt_sizefilt\_size)
@@
$filtered\_disp\leftarrow$ FilterDisp(dispdisp, $K_{mean}$, $K_{sub}$)
@@
return  filtered_depthfiltered\_depth
@@
function FilterDisp(dispdisp, $K_{mean}$, $K_{sub}$)
@@
filledif $led\leftarrow\text{if }den>0\text{ then }\frac{num}{den}\text{ else }mask$ed
return  filledfilled
@@ Appendix C Training Details for Navigation with Reinforcement Learning
Linear Velocity ($v_{t}$) $\pm 0.2$ m/s
Angular Velocity ($\omega_{t}$) $\pm 0.1$ rad/s
Projected Gravity ($n_{t}$) $\pm 0.1$
Goal Position ($p_{t}$) $\pm 0.5$ m, $\pm 0.1$ rad